OpenStudy (anonymous):

can anyone help me get the inverse of this equation... y=2^0.25x-4 +4

OpenStudy (jamesj):

Like any such problem, you have y = f(x). Solve the equation for x and find the function g such that x = g(y). That function g will be your inverse.

OpenStudy (anonymous):

do i do the log since 0.25x-4 is an exponent?

OpenStudy (jamesj):

Yes. You're starting with \[ y = 2^{(x/4) -4} +4 \] hence \[ 2^{(x/4) -4} = y -4 \] Now take the log of both sides.

OpenStudy (anonymous):

i'm stuck at logx-log4=0.25log2-log16

OpenStudy (jamesj):

Starting with my last equation, taking the natural log of both sides, \[ ((x/4) - 4) \ln 2 = \ln (y-4) \] Now, hopefully you can solve that equation for x?